OpenStudy (anonymous):

the sum of two numbers is 100. One number is four more than three times the other number.

OpenStudy (anonymous):

Set up x+y=100 and x=3y+4. Substitute the second equation into the first, solve for y, then solve for x.

OpenStudy (anonymous):

x+(3x+4)=100 4x+4=100 x=96/4=24 second number=3*24+4=76 numbers are 24 and76

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

let us take the first condition,i.e,the sum of two numbers is 100. let the two numbers be x and y. so, x+y=100,satisfies the given condition. now,coming to the second condition,i.e,one number is four more then three times the other number,so,x=3*y-4 satisfies the second condition. substituting x=3*y-4 in the equation,x+y=100,we can solve it and can get the values of x and y as 76 and 24.